OpenStudy (anonymous):

calculate the derivative of 4x/(x^2-36). And find the interval on which it is decreasing. I know that the derivative is (4((x^2)-36)-8x^2)/((x^2)-36)^2 but I don't know how to find where it is decreasing since it doesn't have any critical numbers

OpenStudy (anonymous):

I got the same thing for the derivative, one second on where it is decreasing

OpenStudy (anonymous):

So there are critical numbers at -6 and 6 because they make the denominator 0 which means the fraction will either be 0 or nonexistent--either way they are critical numbers

OpenStudy (anonymous):

You can also find this algebraically by doing the following: 0 = (4((x^2)-36)-8x^2)/((x^2)-36)^2 Multiply by denominator on both sides 0 = 4((x^2)-36)-8x^2 add 8x^2 both sides 8x^2=4((x^2)-36) Divide by 4 both sides 2x^2=x^2 - 36 Subtract x^2 both sides x^2 = 36 Square root both sides x = 6 or x = -6

OpenStudy (anonymous):

So then you can take one derivative from each of the following intervals to find where it's increasing: x<-6 -6<x<6 6<x

OpenStudy (anonymous):

So these are examples f'(-7)=-340/169 f'(0)=-1/9 f'(7)=-340/169 So it is decreasing in all of those intervals because they are all negative derivatives. You can check your answer by looking at the graph of f(x) (attached).

OpenStudy (anonymous):

In that graph you can see that the function is decreasing everywhere, so our answer makes sense!

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Hope I helped! Feel free to ask questions if you have any. Also, wolframalpha.com is a really great resource for performing calculations (not as great at explaining, though).